# Non-Gaussian Simultaneous Autoregressive Models with Missing Data

### v1

@@ 5.2 SEMs with missing data
This section discusses the analysis of the Lucas-1998-HP dataset with missing data in the response variable. We use the age of the house and the response variable of SEMs (logarithm of the house price) as the covariates in the missing data model. The missing data model has three parameters, $\boldsymbol{\psi}=(\psi_{0},\psi_{\textbf{x}^{*}},\psi_{\textbf{y}})$. We set the values $\psi_{0}=0.1$, $\psi_{\textbf{x}^{*}}=0.1$, and $\psi_{\textbf{y}}=-0.1$. With these settings, approximately 50% of the responses are missing, resulting in $n_{u}=2,348$.
Since both $n$ and $n_{u}$ are large, we utilise the HVB-AllB algorithm. We set the number of MCMC iterations $N_{1}$ to 10 and the block size to 10% of $n_{u}$, resulting in 11 blocks. The starting values for the HVB-AllB algorithms are obtained as follows: the variational mean vector ($\boldsymbol{\mu}$) is set to the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau model with only the observed data, $\textbf{y}_{o}$, using MLE. The initial value for $\nu$ is set to 4, and the initial value for $\gamma$ is set to 1. For the parameters of the missing data generating model ($\psi_{0}$, $\psi_{\textbf{x}^{*}}$, and $\psi_{\textbf{y}}$), all coefficients are set to an initial value of 0.01. For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01. The initial values for the missing data ($\mathbf{y}_{u}$) are simulated from the conditional distribution $p(\mathbf{y}_{u}\mid\boldsymbol{\xi}^{(0)},\mathbf{y}_{o})$, where $\boldsymbol{\xi}^{(0)}$ represents the vector of initial parameter values for $\boldsymbol{\xi}$. We use $p=4$ factors.
Similar to the simulation study with missing data, models with Student’s $t$ errors required a relatively large number of iterations to converge. Therefore, we ran the HVB-AllB algorithm for 10,000 iterations for the SEM-Gau and YJ-SEM-Gau models, 40,000 iterations for the SEM-t model, and 20,000 iterations for the YJ-SEM-t model. Convergence plots can be found in Section S9.2 of the online supplement.
@@
The estimated posterior means and 95% credible intervals for selected model parameters across different models are summarised in Table 7, along with the corresponding $\text{DIC}_{5}$ values. According to the $\text{DIC}_{5}$ values, the YJ-SEM-t model is the most appropriate SEM for the Lucas-1998-HP dataset with missing data, followed by the YJ-SEM-Gau model as the second best model.
@@
Table 7: Posterior means and 95% credible intervals for selected parameters of different SEMs for the Lucas-1998-HP dataset with missing data. The table also includes $\text{DIC}_{5}$ values for each model. Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
The estimated parameters from different SEMs are now briefly discussed. For the spatial autocorrelation parameter, $\rho$, the posterior mean estimates across all SEMs are slightly lower than those obtained without missing values, shown in Table 6. However, these estimates still indicate a moderate spatial correlation in the Lucas-1998-HP dataset, consistent with the conclusions drawn from SEMs fitted without missing values in Section 5.1.
The posterior mean estimates for the fixed effects ($\boldsymbol{\beta}$) in the SEMs lead to the same conclusions as those drawn from the Lucas-1998-HP dataset without missing values. For example, the posterior mean estimate of $\beta_{\text{age}}$ in the SEM-Gau model suggests a positive effect of age on house prices, whereas the estimates from the other three models indicate a negative effect. Moreover, the estimated intercepts in the missing data model, ${\psi}_{0}$, are close to the true value of $0.1$ for all models except SEM-t.
The posterior mean estimates of $\nu$ for the SEM-t and YJ-SEM-t models indicate moderate to strong heavy-tailed behavior. Meanwhile, the estimated $\gamma$ values for the YJ-SEM-Gau and YJ-SEM-t models suggest left skewness in the data. These conclusions are consistent with the findings from the analysis without missing data discussed in Section 5.1.
@@ 6 Conclusion
Our article introduces three novel simultaneous autoregressive (SAR) models for non-Gaussian spatial data. Specifically, we extend the conventional spatial error model (SEM-Gau), a widely used SAR model, to accommodate skewed and heavy tailed features of real datasets. The novel SEMs are SEM-t, YJ-SEM-Gau, and YJ-SEM-t. The SEM-t assumes that the error terms follow a Student’s t distribution, allowing heavier tails than a Gaussian distribution. The YJ-SEM-Gau retains Gaussian errors but applies the Yeo-Johnson (YJ) transformation to handle asymmetric response variables. The YJ-SEM-t combines these approaches by assuming Student’s t errors while incorporating the YJ transformation to account for skewness in the response variable. These extensions improve the robustness of SAR models in analysing complex spatial datasets. Additionally, we develop efficient variational Bayes (VB) methods to estimate these models with and without missing data. For cases with missing data, our approach assumes a missing not at random (MNAR) mechanism. Standard VB methods are inadequate for accurately estimating these SEMs in the presence of missing data; therefore, we employ hybrid VB (HVB) algorithms to address this challenge. The posterior distributions obtained from the proposed VB and HVB approximation methods are compared with those of the exact Hamiltonian Monte Carlo (HMC) method.
The empirical results show that: (1) The VB and HVB methods yield posterior density estimates for most model parameters that are similar to those obtained using the HMC method when estimating SEMs with and without missing values; (2) For models with Student’s $t$ errors (SEM-t and YJ-SEM-t), the posterior approximation of the degrees of freedom parameter ($\nu$) from the VB and HVB methods exhibits slight deviations compared to the true value; (3) For both VB and HVB algorithms, models with Student’s t errors require more computation time per iteration than models with Gaussian errors (SEM-Gau and YJ-SEM-Gau); (4) In general, models with Gaussian errors tend to converge in fewer iterations compared to those with Student’s t errors; (5) YJ-SEM-t is able to model complex response variable with skewness and heavy tails and is the best model for the Lucas-1998-HP dataset, both with and without missing data.
@@
This section provides the Yeo and Johnson (YJ) transformation (Yeo and Johnson,, 2000) and its derivatives, which are used to construct the YJ-SEM-Gau and YJ-SEM-t models discussed in Section 2.2 of the main paper.
Let the random variable $y_{i}$ be asymmetric. The Yeo–Johnson (YJ) transformation, denoted $t_{\gamma}(y_{i})$, is used to make $y_{i}$ more symmetric, and is defined by:
@@
where $20<\gamma<$ 2.
The density function of SEMs with YJ transformations includes the derivative of the YJ transformation with respect to $y_{i}$, as shown in Equation (7) of the main paper. This derivative is equal to
@@
Table S1: The definitions of r, M, and the parameters $\boldsymbol{\phi}$ for the four spatial error models. Each SEM has a mean vector of $\mathbf{X}\boldsymbol{\beta}$ and a covariance matrix of the form $\boldsymbol{\Sigma}=\sigma^{2}_{\textbf{e}}\textbf{M}^{-1}$, with $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$.
@@
SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $^{\top}\textbf{A}$ A
SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$ A
YJ-SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e},},\rho,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $^{\top}\textbf{A}$ A
YJ-SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$ A
@@
where the term $p(\textbf{y}\mid\boldsymbol{\xi})$ represents the likelihood of y, and the log-likelihood of y for different SEMs are given in Section 2 of the main paper. We define $p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})=h(\boldsymbol{\xi})$. The terms $h(\boldsymbol{\xi})$ for each SEM, along with the total length of the parameter vector to be estimated in each model (i.e., the length of the vector $\boldsymbol{\xi}$), are provided in Table S2.
Table S2: The term $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})$, and the total number of model parameters $s$ (including the length of the latent vector $\boldsymbol{\tau}$ if exists) for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models. The distribution $p(\boldsymbol{\tau}\mid\boldsymbol{\phi})=\prod_{i=1}^{n}p(\tau_{i}\mid\nu)$ represents the $n$ independent inverse gamma latent variables used in SEMs with Student-$t$ errors (see Equation (2) on the main paper).
@@
Now we explain how variational Bayes (VB) inference is used to approximate the posterior distribution in Equation (S3). We use the variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$, indexed by the variational parameter $\boldsymbol{\lambda}$ to approximate the posterior $p(\boldsymbol{\xi}\mid\textbf{y})$. The VB approach approximates this posterior distribution by minimising the Kullback-Leibler (KL) divergence between $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$and $p(\boldsymbol{\xi}\mid\textbf{y})$. The KL divergence between these two distributions is
@@
[…]
@@
$\mathcal{L}(\boldsymbol{\lambda})=E_{q}\left[\text{log}~{}h(\boldsymbol{\xi})-\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\right],$ (S6)
where $E_{q}\left[\cdot\right]$ denotes the expectation with respect to $q_{\boldsymbol{\lambda}}$.
We employ the Gaussian variational approximation, where we select $q_{\boldsymbol{\lambda}}$ to be a multivariate normal distribution. As a result, the variational parameters $\boldsymbol{\lambda}$ consist of both the mean vector and the distinct elements of the covariance matrix. We impose a factorised covariance structure (Ong et al.,, 2018) on the covariance matrix of $q_{\boldsymbol{\lambda}}$, which reduces the number of distinct elements in the covariance matrix. Under a factor covariance structure for the covariance matrix, the variational distribution is parameterised as $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\sim N(\boldsymbol{\xi};\boldsymbol{\mu},\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})$, where $\boldsymbol{\mu}$ is the $s\times 1$ mean vector, B is an $s\times p$ full rank matrix with $p<<s$, and D is an $s\times s$ diagonal matrix having diagonal elements $\textbf{d}=(d_{1},\ldots,d_{s})$. For all SEMs, the total number of parameters, $s$ is provided in Table S2. We further impose the restriction that the upper triangular elements of B are all zero.
@@
[…]
Minimising the variance of the gradient estimator, $\widehat{{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}}$, is crucial for both stability and fast convergence of the SGA algorithm, which we address using the reparameterisation trick (Kingma and Welling,, 2013; Rezende et al.,, 2014).
We now briefly explain how the reparameterisation trick, combined with a factor covariance structure for the Gaussian variational distribution, is used to obtain efficient gradient estimates in this article. More detailed information can be found in Ong et al., (2018).
$\begin{split}\mathcal{L}(\boldsymbol{\lambda})&=E_{q}\left[\text{log}~{}h(\boldsymbol{\xi})-\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\right]\\ &=E_{f_{\boldsymbol{\zeta}}}\left[\text{log}~{}h(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))-\text{log}~{}q_{\boldsymbol{\lambda}}(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))\right],\end{split}$ (S8)
$\begin{split}\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\zeta}}}\left[\nabla_{\boldsymbol{\lambda}}~{}\text{log}~{}h(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))-\nabla_{\boldsymbol{\lambda}}~{}\text{log}~{}q_{\boldsymbol{\lambda}}(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))\right],\\ &=E_{f_{\boldsymbol{\zeta}}}\left[\frac{du(\boldsymbol{\zeta,\boldsymbol{\lambda}})^{\top}}{d\boldsymbol{\lambda}}\{\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\xi})-\nabla_{\boldsymbol{\xi}}\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\}\right],\\ \end{split}$ (S9)
@@
Algorithm S1 outlines the VB algorithm for estimating SEMs with full data. The unbiased estimate of the gradient, $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}$, presented in step 4 of Algorithm S1 can be constructed using the reparameterisation trick described earlier, with a single sample drawn from $f_{\boldsymbol{\zeta}}(\boldsymbol{\zeta})$.
@@
[…]

### v2

@@ 8.2 SEMs with missing responses
This section analyses the Lucas-1998-HP dataset with missing values in the response variable (logarithm of the house price). Starting from the complete dataset, missing responses are generated using the logistic regression model in Equation (11), with the age of the house and the response variable y (from the SEMs) serving as covariates. The missing data model parameters are set to $\psi_{0}=0.1$, $\psi_{\textbf{x}^{*}}=0.1$, and $\psi_{\textbf{y}}=-0.1$. This configuration results in approximately 50% of the responses being missing, yielding $n_{u}=2{,}348$.
[…]
@@
Table 7: Posterior means and 95% credible intervals for selected parameters of different SEMs for the Lucas-1998-HP dataset with missing data. The table also includes $\text{DIC}_{5}$ values for each model. Parameters labelled as ’NA’ indicate that they are not applicable to the corresponding model.
@@
The estimated parameters from different SEMs are now briefly discussed. For the spatial autocorrelation parameter, $\rho$, the posterior mean estimates across all SEMs are slightly lower than those obtained using the full dataset without missing values, except under the SEM-t (see Table 6). However, these estimates still indicate a moderate spatial correlation in the Lucas-1998-HP dataset, consistent with the conclusions drawn from SEMs fitted without missing values in Section 8.1. The posterior mean estimates for the fixed effects ($\boldsymbol{\beta}$) in the SEMs lead to the same conclusions as those drawn from the Lucas-1998-HP dataset without missing values. For example, the posterior mean estimate of $\beta_{\text{age}}$ in the SEM-Gau suggests a positive effect of age on house prices, whereas the estimates from the other three models indicate a negative effect. Moreover, the estimated posterior means of intercept in the missing data model, ${\psi}_{0}$, are close to the true value of $0.1$ for all models except SEM-t. The posterior mean estimates of $\nu$ for the SEM-t and YJ-SEM-t indicate moderate to strong heavy-tailed behaviour. Similarly, the estimated posterior mean values of $\gamma$ for the YJ-SEM-Gau and YJ-SEM-t suggest left skewness in the response. These conclusions are consistent with the findings from the analysis without missing data discussed in Section 8.1.
@@ 9 Conclusion
Our article introduces three novel simultaneous autoregressive (SAR) models for non-Gaussian spatial data. Specifically, we extend the conventional spatial error model (SEM-Gau), a widely used SAR model that assumes Gaussian errors, to accommodate skewed and heavy-tailed features of real datasets. The novel SEMs are SEM-t, YJ-SEM-Gau, and YJ-SEM-t. The SEM-t assumes that the error terms follow a Student’s $t$ distribution, allowing for heavier tails than the Gaussian case. The YJ-SEM-Gau retains Gaussian errors but applies the Yeo–Johnson (YJ) transformation to handle asymmetric response variables. The YJ-SEM-t combines these approaches by assuming Student’s $t$ errors while incorporating the YJ transformation. These extensions improve the robustness of SAR models for analysing non-Gaussian spatial datasets. Additionally, we develop efficient variational Bayes (VB) methods to estimate these models with and without missing data. In cases with missing data, we assume that the responses are missing under the missing not at random (MNAR) mechanism. Standard VB methods are inadequate for accurately estimating these SEMs in the presence of missing data; therefore, we employ hybrid VB (HVB) algorithms to address this challenge. The posterior distributions obtained from the proposed VB and HVB methods are compared with those of the exact Hamiltonian Monte Carlo (HMC) method, demonstrating both accuracy and substantial computational gains. Although the proposed model extensions are illustrated using the SEM, both the modelling framework and the associated VB/HVB inference methods are general and can be readily adapted to other SAR-type models.
The empirical results show that: (1) The VB and HVB methods yield posterior density estimates for most model parameters that are similar to those obtained using the HMC method when estimating SEMs with and without missing values; (2) For models with Student’s $t$ errors (SEM-t and YJ-SEM-t), the posterior approximation of the degrees of freedom parameter ($\nu$) from the VB and HVB methods exhibits slight deviations compared to the true value; (3) For both VB and HVB algorithms, models with Student’s $t$ errors require more computation time per iteration than models with Gaussian errors (SEM-Gau and YJ-SEM-Gau); (4) In general, models with Gaussian errors tend to converge in fewer iterations compared to those with Student’s $t$ errors; (5) YJ-SEM-t is able to model complex response variables with skewness and heavy tails, and it provides the best fit to the Lucas-1998-HP dataset, both with and without missing data.
Future work could explore more flexible transformation families, such as Tukey’s g-and-h transformation (Tukey,, 1977; Martinez and Iglewicz,, 1984), which jointly accommodate skewness and heavy tails. In addition, while the proposed non-Gaussian SEMs are developed for univariate spatial responses, extending the framework to multivariate and spatio-temporal settings would enable joint modelling of multiple non-Gaussian variables and dynamic processes. These extensions would substantially broaden the applicability of the models to a wider range of real-world spatial and spatio-temporal datasets.
@@
This section provides the Yeo and Johnson (YJ) transformation (Yeo and Johnson,, 2000) and its derivatives, which are used to construct the YJ-SEM-Gau and YJ-SEM-t discussed in Section 2.2 of the main paper.
Let the random variable $y_{i}$ be asymmetric. The YJ transformation, denoted $t_{\gamma}(y_{i})$, is often applied to reduce asymmetry and make $y_{i}$ more symmetric. It is defined as:
@@
where $0<\gamma<2$.
The density function of SEMs with YJ transformations includes the derivative of the YJ transformation with respect to $y_{i}$, as shown in Equation (8) of the main paper. This derivative is given by:
@@
In this section, we present the variational Bayes (VB) algorithm for estimating the SEMs without missing values introduced in Section 2 of the main paper.
Table S1: The definitions of r, M, and the parameters $\boldsymbol{\phi}$ for SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t. For SEM-Gau and SEM-t, the density of the response vector conditional on $\boldsymbol{\xi}$ follows a multivariate Gaussian distribution with mean $\textbf{X}\boldsymbol{\beta}$ and covariance matrix $\sigma^{2}_{\textbf{e}}\textbf{M}^{-1}$, where $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$.
@@
SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\textbf{A}$
SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$
YJ-SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e},},\rho,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\textbf{A}$
YJ-SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$
@@
where the term $p(\textbf{y}\mid\boldsymbol{\xi})$ represents the likelihood, and the log-likelihood for different SEMs are given in Section 2 of the main paper. We define $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})$. The terms $h(\boldsymbol{\xi})$ for each SEM, along with the total length of the parameter vector to be estimated in each model (i.e., the length of the vector $\boldsymbol{\xi}$), are provided in Table S2.
Table S2: The term $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})$, and the total number of model parameters $s$ (including the length of the latent vector $\boldsymbol{\tau}$ if exists) for the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t. The density $p(\boldsymbol{\tau}\mid\boldsymbol{\phi})=\prod_{i=1}^{n}p(\tau_{i}\mid\nu)$ represents the $n$ independent inverse gamma latent variables used in SEMs with Student-$t$ errors (see Equation (2) of the main paper).
@@
Now we explain how VB inference is used to approximate the posterior distribution in Equation (S3). We use the variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$, indexed by the variational parameter $\boldsymbol{\lambda}$ to approximate the posterior $p(\boldsymbol{\xi}\mid\textbf{y})$. The VB approach approximates this posterior distribution by minimising the Kullback-Leibler (KL) divergence between $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$ and $p(\boldsymbol{\xi}\mid\textbf{y})$. The KL divergence between these two distributions is
@@
Minimising KL divergence between $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$ and $p(\boldsymbol{\xi}\mid\textbf{y})$ is equivalent to maximising evidence lower bound (ELBO) on the marginal log-likelihood, $\text{log}~p(\textbf{y})$, denoted by $\mathcal{L}(\boldsymbol{\lambda})$, with $p(\textbf{y})=\int~p(\textbf{y}\mid\boldsymbol{\xi})p\left(\boldsymbol{\xi}\right)d\boldsymbol{\xi}$ (Blei et al.,, 2017). The ELBO is
@@
$\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{q}\left[\text{log}~h(\boldsymbol{\xi})-\text{log}~q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\right],$ (S6)
where $\mathbb{E}_{q}\left[\cdot\right]$ denotes the expectation with respect to $q_{\boldsymbol{\lambda}}$.
We employ the Gaussian variational approximation, where we select $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$ to be a multivariate Gaussian distribution. As a result, the variational parameters $\boldsymbol{\lambda}$ consist of both the mean vector and the distinct elements of the covariance matrix. We impose a factor covariance structure (Ong et al.,, 2018) on the covariance matrix of $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$, which reduces the number of distinct elements in the covariance matrix. Under a factor covariance structure for the covariance matrix, the variational distribution is parameterised as $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\sim N(\boldsymbol{\xi};\boldsymbol{\mu},\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})$, where $\boldsymbol{\mu}$ is the $s\times 1$ mean vector, B is an $s\times p$ full rank matrix with $p<<s$, and D is an $s\times s$ diagonal matrix with positive diagonal elements $\textbf{d}=(d_{1},\ldots,d_{s})$. For all SEMs, the total number of parameters, $s$ is provided in Table S2. We further impose the restriction that the upper triangular elements of B are all zero.
@@
$\boldsymbol{\lambda}^{(t+1)}=\boldsymbol{\lambda}^{(t)}+\mathbcal{a}^{(t)}\circ\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$, (S7)
and continues until a stopping criterion is satisfied, where $\mathbcal{a}^{(t)}$ is a vector of element-wise learning rates at iteration $t$. The individual learning rates, denoted by $\mathcal{a}^{(t)}_{i}$, are typically chosen to satisfy the Robbins–Monro conditions: $\sum_{t}\mathcal{a}^{(t)}_{i}$$\infty$ and $\sum_{t}(\mathcal{a}^{(t)}_{i})^{2}<\infty$, which ensure convergence of the sequence $\boldsymbol{\lambda}^{(t)}$ to a local optimum as $t\rightarrow\infty$, under regularity conditions (Robbins and Monro,, 1951; Bottou,, 2010). The symbol $\circ$ denotes the element-wise product of two vectors. To speed up convergence, we use the ADADELTA algorithm for adaptive learning rates (Zeiler,, 2012), detailed in Section S3. Further, minimising the variance of the gradient estimator, $\widehat{{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}}$, is crucial for both stability and fast convergence of the SGA algorithm, which we address using the reparameterisation trick (Kingma and Welling,, 2013; Rezende et al.,, 2014).
We now briefly explain how the reparameterisation trick, combined with a factor covariance structure for the Gaussian variational distribution, is used to obtain efficient gradient estimates in this paper. More detailed information can be found in Ong et al., (2018).
$\begin{split}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{q}\left[\text{log}~h(\boldsymbol{\xi})-\text{log}~q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\right]\\ &=\mathbb{E}_{f_{\boldsymbol{\zeta}}}\left[\text{log}~h(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))-\text{log}~q_{\boldsymbol{\lambda}}(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))\right],\end{split}$ (S8)
@@
Algorithm S1 outlines the VB algorithm for estimating SEMs with full data. The unbiased estimate of the gradient, $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}$, in step 4 of Algorithm S1 can be constructed using the reparameterisation trick described earlier, with a single sample drawn from $f_{\boldsymbol{\zeta}}(\boldsymbol{\zeta})$.
@@
5:  Set adaptive learning rates for the variational means, $\mathbcal{a}^{(t)}_{\boldsymbol{\mu}}$ and the variational parameters $\textrm{vech}(\textbf{B})$ and d, $\mathbcal{a}_{\textrm{vech}{(\textbf{B})}}^{(t)}$ and $\mathbcal{a}_{\textbf{d}}^{(t)}$, respectively, using ADADELTA described in Section S3 of the online supplement.
6:  Set $\boldsymbol{\mu}^{(t+1)}=\boldsymbol{\mu}^{(t)}+\mathbcal{a}^{(t)}_{\boldsymbol{\mu}}\circ\widehat{\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$.
7:  Set $\textrm{vech}(\textbf{B})^{(t+1)}=\textrm{vech}(\textbf{B})^{(t)}+\mathbcal{a}_{\text{vech}(\textbf{B})}^{(t)}\circ\widehat{\nabla_{\textrm{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$.
8:  Set $\textbf{d}^{(t+1)}=\textbf{d}^{(t)}+\mathbcal{a}_{\textbf{d}}^{(t)}\circ\widehat{\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$.